# GMG: A Video Prediction Method Based on Global Focus and Motion Guided

### v2

@@ III-C Motion Guided Module
$\bm{u_{t}}=\sigma(\texttt{Conv2d}(\texttt{Concat}([\bm{\hat{{H_{t}^{g}}}},\mathcal{F}_{t-1}]))),$ (7)
$\bm{r_{t}}=\sigma(\texttt{Conv2d}(\texttt{Concat}([\bm{\hat{{H_{t}^{g}}}},\mathcal{F}_{t-1}]))),$ (8)
$\bm{z_{t}}=\tanh(\texttt{Conv2d}(\texttt{Concat}([\bm{\hat{{H_{t}^{g}}}},\mathcal{F}_{t-1}\circ\bm{r_{t}}]))),$ (9)
where $\bm{u_{t}}$ is the Update Gate, $\bm{r_{t}}$ is the Reset Gate, and $\bm{z_{t}}$ is the reset feature of the current moment. In this equation, $\bm{\hat{{H_{t}^{g}}}}$ will be processed by a stride-2 convolution into a shape of : $\bm{\hat{{H_{t}^{g}}}}\in\mathbb{R}^{T\times\frac{C}{4}\times\frac{H}{2}\times\frac{W}{2}}$.
@@
$\bm{G_{T}}=\mathcal{AAP}(\bm{\hat{{H_{t}^{g}}}}).$ (11)
Based on this dynamic feature, we calculate the dynamic balance factor $\alpha_{d}$ and dynamic decay factor $\beta_{d}$:
@@
Next, we balance the dynamics and the default $\bm{\alpha}$ and $\bm{\beta}$:
$\bm{\alpha}=\mathscr{T}(\bm{\alpha_{d}},\bm{\alpha^{\prime}}),$ (13)
$\bm{\beta}=\mathscr{T}(\bm{\beta_{d}},\bm{\beta^{\prime}}),$ (14)
where, $\bm{\alpha^{\prime}}$ and $\bm{\beta^{\prime}}$ are the initial parameters, respectively. $\mathscr{T}$ is the transfer fuction. In this paper, the transfer function is used to fuse the values of $\bm{\alpha}$ and $\bm{\beta}$ by taking the average of the two inputs.
We also introduce an exponential decay mechanism:
$\bm{D_{e}}=\texttt{exp}(-\bm{\beta}\cdot\texttt{TimeSteps}),$ (15)
where TimeSteps is a tensor with the same time steps as $\mathcal{D}$, TimeSteps $\in$ $\mathbb{R}^{T\times 1\times 1\times 1}$. This formula used to simulate the effect of exponential decay, meaning that the further in time, the smaller the weight.
@@
We use DeformableConv [25] to better adapt the motion variations. In addition, we propose two GMG variants to adapt to different tasks (GMG-s, GMG-m, reffer to Fig. 3). To accommodate large-sample datasets, a simplified GFM is provided to reduce computational complexity:
@@
III-D Overall structure for GMG
@@
$i_{t}^{\prime}=\sigma\left(W_{xi^{\prime}}*X_{t}+W_{hi^{\prime}}*{\hat{M}}_{t-1}+b_{i}^{\prime}\right)$ (31)
@@
$M_{t}=f_{t}^{\prime}\circ{\hat{M}}_{t-1}+i_{t}^{\prime}\circ{\hat{C}}_{t}^{\prime}$ (33)
@@
Figure 3: Diagram of GMG and Its Variants’ Architecture
@@ IV-A Datasets and Metrics
In this paper, we conduct experiments on five datasets: CIKM2017[26], Shanghai2020[27], Taxibj[28], WeatherBench[29] and Moving MNIST[31].
@@
Implementation Details: The GMG model is trained using an $L1+L2$ loss to simultaneously enhance the sharpness and smoothness of the generated frames. The specific model parameters, hyperparameters (such as batch size, learning rate, and training epochs), as well as the training hardware for each dataset, are detailed in Table I.
Metrics: We use common metrics to evaluate the predictive performance of the model, including MSE, MAE, RMSE, PSNR and SSIM. Specifically, for rainfall data, we also include the Critical Success Index (CSI) to assess forecast accuracy. A CSI value closer to 1 indicates higher prediction accuracy.
TABLE I: Experimental settings for different datasets.
@@
Moving MNIST 4/16 (1×64×64) GMG-L $10\rightarrow 10$ 600 0.0003 32 1 * TESLA V100
@@ IV-B Quantitative Evaluation
To analyze the performance of GMG compared to other SOTA models, we summarize the experimental results. Tables II, III, IV, and V present these quantitative results. Numbers highlighted in Red Bold indicate the best performance for the corresponding metric, while numbers marked with a Blue Underlined represent the second-best performance. In this study, we select deterministic models for comparison, including CNN, RNN, CNN-RNN, and ViT-based models. The selected models also include mainstream SOTA models such as SimVP and its variants, SwinLSTM, and WaST.
Table II presents the performance of baseline models and GMG on the CIKM2017 and Shanghai2020 datasets. GMG achieves leading performance across all metrics. Compared to the best-performing baseline models, GMG reduces the MSE by 8% on CIKM2017 and 4.1% on Shanghai2020. Additionally, SSIM improves by 0.0107 on CIKM2017 and 0.0012 on Shanghai2020. Notably, RNN-based models (e.g., MIM, MotionRNN, PredRNN-V2) tend to perform better in these two rainfall datasets. This may be attributed to the superior temporal modeling capability, spatial information capture, and effective long-term dependency modeling of recurrent neural networks.
Table III compares the performance of different models on the TaxiBJ dataset. Compared to the current best model, WaST, GMG achieves a new SOTA in MAE, SSIM, and PSNR. Compared to the baseline model MotionRNN, GMG reduces MSE by 20.63% and lowers MAE by approximately 7.9%.
Tables IV and V present the experimental results on the WeatherBench and Moving MNIST datasets, respectively, where GMG achieves comprehensive superiority across all evaluation metrics. With minor modifications, GMG, as an RNN-based model, outperforms the currently popular CNN and ViT models across different tasks. This provides a new perspective on utilizing recurrent neural networks for complex video prediction tasks in the future.
TABLE II: Quantitative results on CIKM2017 and Shanghai2020 datasets.
@@
TABLE III: Quantitative results in Taxibj datasets.
@@
TABLE IV: Quantitative results in WeatherBench datasets.
@@
TABLE V: Quantitative results in Moving MNIST datasets.
@@ IV-C Qualitative Evaluation
Figure 4: Visualizations on CIKM2017 (upper) and Shanghai2020 (lower). The time labeled in the figure represents the input data time/predicted time.
@@
Fig. 5 presents the experimental results on Taxibj, comparing the outcomes of TAU and WaST. In the main regions where errors occur (marked by green and red boxes), GMG consistently maintains the lowest error. This further demonstrates the effectiveness of the Global Focus Model in accurately capturing long-range dependencies.

### v3

@@ III-E Motion Guided Module
$\bm{u_{t}}=\sigma(\texttt{Conv2d}_{W_{\mathcal{\bm{U}}}}(\texttt{Concat}([\bm{\hat{{H_{t}^{g}}}},\mathcal{F}_{t-1}]))),$ (19)
$\bm{r_{t}}=\sigma(\texttt{Conv2d}_{W_{\mathcal{\bm{R}}}}(\texttt{Concat}([\bm{\hat{{H_{t}^{g}}}},\mathcal{F}_{t-1}]))),$ (20)
$\bm{z_{t}}=\tanh(\texttt{Conv2d}_{W_{\mathcal{\bm{Z}}}}(\texttt{Concat}([\bm{\hat{{H_{t}^{g}}}},\mathcal{F}_{t-1}\circ\bm{r_{t}}]))),$ (21)
where $\bm{u_{t}}$ is the Update Gate, $\bm{r_{t}}$ is the Reset Gate, ${W_{\mathcal{\bm{U}}}},{W_{\mathcal{\bm{R}}}},{W_{\mathcal{\bm{Z}}}}$ represent convolution kernel, and $\bm{z_{t}}$ is the reset feature of the current moment. In this equation, $\bm{\hat{{H_{t}^{g}}}}$ will be processed by a stride-2 convolution into a shape of : $\bm{\hat{{H_{t}^{g}}}}\in\mathbb{R}^{T\times\frac{C}{4}\times\frac{H}{2}\times\frac{W}{2}}$.
@@
[…]
Based on this dynamic feature, we calculate the dynamic balance factor $\alpha_{d}$ and dynamic decay factor $\beta_{d}$ through a lightweight linear projection:
@@
These two factors are designed to model the intrinsic physical processes of non-rigid motion. Specifically, $\bm{\alpha}$ reflects the motion “growth tendency”, accounting for momentum preservation and internal deformation, while $\bm{\beta}$ captures the “dissipative decay”, simulating energy loss over time due to friction or structural damping. Instead of assigning them as fixed hyperparameters, we adopt a data-driven approach that combines dynamic estimates $\bm{\alpha_{d}},\bm{\beta_{d}}$ with learnable priors $\bm{\alpha^{\prime}},\bm{\beta^{\prime}}$, using an element-wise arithmetic mean to ensure robustness:
$\bm{\alpha}=\mathscr{T}(\bm{\alpha_{d}},\bm{\alpha^{\prime}}),\quad\bm{\beta}=\mathscr{T}(\bm{\beta_{d}},\bm{\beta^{\prime}}).$ (25)
where $\bm{\alpha^{\prime}}$ and $\bm{\beta^{\prime}}$ are the initial parameters, respectively. $\mathscr{T}(\cdot)$ is the transfer function. In this paper, the transfer function $\mathscr{T}(\cdot)$ fuses the two inputs by computing their element-wise arithmetic mean. This fusion enables adaptive learning while preserving stable physical interpretability across different input distributions.
To further reflect the temporal attenuation of motion influence, we introduce a Time Delay mechanism:
$\bm{D_{e}}=\exp(-\bm{\beta}\cdot\texttt{TimeSteps}),$ (26)
where TimeSteps is a monotonically increasing time index tensor, $\texttt{TimeSteps}\in\mathbb{R}^{T\times 1\times 1\times 1}$. The exponential decay formulation faithfully reflects physical systems where motion gradually weakens with time, such as damping oscillation or deformation relaxation. This allows the model to automatically prioritize recent motion cues and suppress the influence of outdated dynamics.
@@
We use DeformableConv [29] to better adapt the motion variations. Compared with MotionGRU, MGM offers two key advantages. First, it explicitly models non-rigid deformations, enabling the capture of growth, dissipation, and morphological changes of targets during motion. Second, its factors have clear physical interpretations; combined with data-driven dynamic estimation and learnable priors, this allows the model to remain robust under varying input distributions and to effectively balance near-term and long-term information in video prediction over extended time horizons, thereby substantially improving its ability to capture complex dynamic targets and enhancing generalization performance.
In addition, we propose two GMG variants to adapt to different tasks (GMG-s, GMG-m, refer to Fig. 3). To accommodate large-sample datasets, a simplified GFM is provided to reduce computational complexity:
@@
To evaluate the efficiency of key modules, we compare their computational complexity and parameter sizes (see Table I). The GFM module has a time complexity of $\mathcal{O}(BCHW)$ and a low memory footprint of $\mathcal{O}(BC)$, offering effective global modeling with high efficiency. In contrast, the MGM module is even more lightweight, with minimal computational and memory cost, making it well-suited for capturing fine-grained local motion.
In our model variants, GMG-s adopts a simplified version of GFM, while GMG-m removes the computationally expensive SAM module but retains GFM and MGM, achieving a good balance between performance and efficiency.
TABLE I: Module complexity and size comparison.
Module Time Space Params FLOPs
SAM $\mathcal{O}(BH^{2}W^{2})$ $\mathcal{O}(BH^{2}W^{2})$ 1.08M 7.17G
GFM $\mathcal{O}(BCHW)$ $\mathcal{O}(BC)$ 0.53M 2.25G
MGM $\mathcal{O}(BCHW)$ $\mathcal{O}(BCHW)$ 0.10M 0.13G
III-F Overall structure for GMG
@@
$i_{t}^{\prime}=\sigma\left(W_{xi^{\prime}}*X_{t}+W_{hi^{\prime}}*{{M}}_{t-1}+b_{i}^{\prime}\right)$ (42)
@@
$M_{t}=f_{t}^{\prime}\circ{{M}}_{t-1}+i_{t}^{\prime}\circ{\hat{C}}_{t}^{\prime}$ (44)
@@
Figure 3: Diagram of GMG and Its Variants’ Architecture.
@@ IV-A Datasets and Metrics
In this paper, we conduct experiments on six datasets: CIKM2017[30], Shanghai2020[31], Taxibj[32], WeatherBench[33], Moving MNIST[35] and Typhoon[36].
@@
Typhoon: The Digital Typhoon Dataset is a satellite image dataset designed for machine learning research on tropical cyclones. We randomly selected 10,140 samples from the Western Pacific (WP) and Around Australia (AU) regions as the training set, and 1,000 samples as the test set. Each sample consists of four input images and four output images, where each image has a spatial resolution of $128\times 128$ pixels and corresponds to a 1-hour interval in time. Compared with other datasets used in this study, this dataset further evaluates the model’s ability to predict fine-grained details.
Implementation Details: To ensure stable training and smooth predictions, the GMG model is optimized using $L_{2}$ loss function, which penalizes larger errors more heavily and corresponds to the assumption of Gaussian-distributed noise. The specific model parameters, hyperparameters (Such as Batch size, Learning rate, Attention hidden layers and Training epochs), as well as the training hardware for each dataset, are detailed in Table II.
Metrics: We use common metrics to evaluate the predictive performance of the model, including MSE, MAE, RMSE, PSNR and SSIM.
Specifically, for rainfall data, we also include the Critical Success Index (CSI) to assess forecast accuracy:
$\text{CSI}=\frac{TP}{TP+FP+FN}$ (47)
where $TP$ represents true positives (correctly predicted positive events), $FP$ denotes false positives (events predicted as positive but actually negative), and $FN$ stands for false negatives (events that were not predicted but are actually positive). A higher CSI value indicates better model performance, with a maximum possible score of 1. CSI is widely used in applications such as weather forecasting, event prediction, and video anomaly detection.
TABLE II: Experimental settings for different datasets.
@@
Moving MNIST 4/16 (1×64×64) GMG-L $10\rightarrow 10$ 600/2000 0.0003 32 1 * TESLA V100
Typhoon 4/16 (1×128×128) GMG-L $4\rightarrow 4$ 100 0.0003 64 1 * TESLA V100
@@ IV-B Quantitative Evaluation
To analyze the performance of GMG compared to other SOTA models, we summarize the experimental results. Tables III, IV, V, and VI present these quantitative results. Numbers highlighted in Red Bold indicate the best performance for the corresponding metric, while numbers marked with a Blue Underlined represent the second-best performance. In this study, we select deterministic models for comparison, including CNN, RNN, CNN-RNN, and ViT-based models. The selected models also include mainstream SOTA models such as SimVP and its variants, SwinLSTM, and WaST.
Table III presents the performance of baseline models and GMG on the CIKM2017 and Shanghai2020 datasets. GMG achieves leading performance across all metrics. Compared to the best-performing baseline models, GMG reduces the MSE by 8% on CIKM2017 (Compare with MIM) and 4.1% on Shanghai2020 (Compare with PredRNN-V2). Additionally, SSIM improves by 0.0107 on CIKM2017 and 0.0012 on Shanghai2020. Notably, RNN-based models (e.g., MIM, MotionRNN, PredRNN-V2) tend to perform better in these two rainfall datasets. This may be attributed to the superior temporal modeling capability, spatial information capture, and effective long-term dependency modeling of recurrent neural networks.
Table IV compares the performance of different models on the TaxiBJ dataset. Compared to the current best model, WaST, GMG achieves a new SOTA in MAE, SSIM, and PSNR. Compared to the baseline model MotionRNN, GMG reduces MSE by 20.63% and lowers MAE by approximately 7.9%.
Tables V and VI present the experimental results on the WeatherBench and Moving MNIST datasets, respectively, where GMG achieves comprehensive superiority across all evaluation metrics. With minor modifications, GMG, as an RNN-based model, outperforms the currently popular CNN and ViT models across different tasks. This provides a new perspective on utilizing recurrent neural networks for complex video prediction tasks in the future.
Specifically, on the Moving MNIST dataset, we trained the GMG model for 2000 epochs, and the results show a significant performance improvement. Compared to GMG trained for a standard (shorter) number of epochs, the longer-trained GMG achieved noticeable gains across all evaluation metrics. This further demonstrates the stability and generalization capability of the GMG architecture under extended training, providing strong support for further optimization of complex video prediction tasks.
Table VII presents the quantitative results on the Typhoon dataset. GMG achieves the best performance across all evaluation metrics. Compared to the MotionRNN, GMG reduces the MSE by 1.2% and reduced RMSE from 12.67 to 12.60. Furthermore, GMG surpasses WaST in both MSE and MAE. Notably, Typhoon is a higher-resolution dataset with complex and chaotic motion patterns, yet GMG maintains its leading performance, demonstrating strong generalization and robustness in challenging real-world scenarios.
TABLE III: Quantitative results on CIKM2017 and Shanghai2020 datasets.
@@
TABLE IV: Quantitative results in Taxibj datasets.
@@
TABLE V: Quantitative results in WeatherBench datasets.
@@
TABLE VI: Quantitative results in Moving MNIST datasets. (GMG* denotes the results obtained after 2000 training epochs.)
@@
GMG* (Ours) 14.2479 48.1517 0.9700 26.3387
TABLE VII: Quantitative results in Typhoon datasets.
Model MSE (↓) MAE/10 (↓) RMSE (↓)
ConvLSTM[25] 224.389 125.000 14.9796
PredRNN[5] 216.077 122.643 14.6995
PredRNN++[6] 183.307 113.522 13.539
MIM[7] 222.208 124.240 14.9066
MotionRNN[8] 160.656 107.228 12.6750
MAU[23] 166.581 109.660 12.9066
PredRNN-V2[9] 183.416 114.336 13.5431
SimVP-gSTA[12] 204.329 119.063 14.2943
TAU[13] 197.883 116.552 14.0671
WaST[34] 168.903 106.952 12.9962
GMG (Ours) 158.794 106.381 12.6013
@@ IV-C Qualitative Evaluation
Figure 4: Visualizations on CIKM2017 (upper) and Shanghai2020 (lower). “+min” indicates the forecast lead time of the predicted images, while “-min” represents the time of the input images.
Fig. 5 presents the error analysis for a typical case from the CIKM 2017 dataset. Compared to the Shanghai 2020 dataset, this dataset is more challenging to predict, yet GMG, through its ability to adaptively model the dynamic evolution of non-rigid meteorological patterns, is able to effectively correct predictions in regions with larger errors. Despite having limited input information, GMG is still able to make accurate predictions for rainfall areas, capturing the changes in precipitation processes and the movement of clouds, which are complex weather phenomena.
Figure 5: Error visualizations on CIKM2017 dataset, “GMG Pred.” indicate the real prediction results made by GMG.
Fig. 6 presents the error analysis for a typical case from Shanghai2020. Through visual comparison with the green box, it is evident that the GMG model exhibits smaller errors in more complex regions compared to other models. The prediction results from GMG show clear details with smooth edge transitions. Compared to traditional CNN or RNN models, GMG demonstrates stronger capability in detail reconstruction in these critical areas, resulting in smaller errors.
Figure 6: Error visualizations on Shanghai2020 dataset, “GMG Pred.” indicate the real prediction results made by GMG.
@@
Fig. 7 presents the experimental results on Taxibj, comparing the outcomes of TAU and WaST. In the main regions where errors occur (marked by green and red boxes), GMG consistently maintains the lowest error. This further demonstrates the effectiveness of the Global Focus Module in accurately capturing long-range dependencies.
Figure 7: Visualizations on Taxibj, Error = $\left|\text{Prediction}-\text{Target}\right|$ , we amplify the error for better comparison.
Fig. 8 presents the visualization results of GMG on the WeatherBench and Moving MNIST datasets. In the WeatherBench section, we display the ground truth temperature maps, GMG’s predictions, and the corresponding error maps in comparison with other models such as TAU and WasT. We specifically highlight the main error-prone regions over the equatorial Pacific area with red boxes. It is clearly observed that GMG achieves the smallest prediction errors in these sensitive regions, with error colors tending toward green (indicating lower error), while other models exhibit more significant deviations. This demonstrates that even though temperature variations in WeatherBench are relatively gradual, GMG leveraging the Global Focus Module and Motion Guided Module can still accurately capture subtle temperature changes, showcasing strong modeling capability.
In the Moving MNIST section, we compare the performance of GMG at different training stages (600 epochs vs. 2000 epochs, the latter denoted as GMG*) and analyze two representative examples: one involving the digit “8” and the other the digit “6”.
In the case of digit “8”, we observe that the TAU model generates blurry and structurally broken predictions in later time steps ($T=8\sim 10$), especially failing to maintain the closed loop of the central ring (as highlighted by red boxes). In contrast, GMG trained with 600 epochs is already able to preserve the overall shape and loop closure of the digit “8”, though some edge blur and detail loss remain. GMG* (trained for 2000 epochs), however, produces much sharper and more coherent predictions, with clear boundaries and complete structure—almost indistinguishable from the ground truth. This indicates that longer training significantly enhances the model’s ability to capture fine-grained details and improves prediction stability.
In the case of digit “6”, TAU’s predictions suffer from deformation and displacement, particularly around the lower curved region of the digit. GMG (600 epochs) performs better, preserving the general outline of the digit “6”, although the stroke appears slightly thin and lacks fullness. GMG* further improves this by accurately restoring the shape, maintaining consistent thickness and spatial continuity (as shown by green boxes), suggesting the model’s strong ability to track rigid object contours and generate more natural and stable predictions.
Figure 8: Visualizations on WeatherBench (upper) and Moving MNIST datasets (lower) , bias = Prediction - Target, which can reflect the extent to which the model tends to predict colder or warmer temperatures.
Fig. 9 presents the visualization results for Typhoon. The visualization results show that GMG performs exceptionally well on datasets with high resolution and complex dynamic patterns, such as typhoons. Compared to WaST, GMG exhibits lower and more localized prediction errors in key regions like the typhoon eye and strong convective bands, preserving spatial details more effectively. Additionally, GMG produces more temporally coherent forecasts across consecutive time steps, demonstrating stronger sequence modeling capabilities, making it well-suited for fine-grained prediction of complex weather systems.
From the above visualization results, we observe that GMG can adapt to various types of video prediction tasks, regardless of whether the predicted subject is a rigid or non-rigid body. Moreover, it achieves superior performance in terms of prediction accuracy and object detail preservation. This demonstrates that GMG can serve as a general-purpose video prediction model, making it applicable to real-world complex video prediction tasks.
Figure 9: Visualizations on Typhoon, Error = $\left|\text{Prediction}-\text{Target}\right|$ , we amplify the error for better comparison.
IV-D Ablation Study
To validate the contributions of GMG’s individual components, we conduct ablation experiments on the CIKM2017 dataset, with results presented in Table VIII. In addition to controlled experiments on key components—GFM, MGM, and SAM—we include comparisons with baseline models, PredRNN and MotionRNN. The results show that adding GFM, SAM, or MGM improves predictive performance over the baselines. Furthermore, compared to MotionGRU, model (A), which incorporates the MGM module to account for growth or dissipation processes, better captures rainfall data characteristics. By comparing Model (A), (B), and (C), MGM emerges as the most impactful component in enhancing baseline performance, followed by GFM, further validating the effectiveness of the proposed model design. Model (D) incorporates both the GFM and SAM modules, and even without using MGM, it achieves a significant performance improvement, indicating the positive roles of GFM and SAM in feature extraction and spatial modeling. Additionally, the lightweight version, GMG-s, achieves competitive results, offering a viable pathway for future lightweight modeling efforts. We also provide the parameter count and training time for each model, offering a reference for practical model deployment.
To compare the impact of each module on training and inference efficiency, we report the parameter count (Params), computational complexity (FLOPs), and inference speed (FPS, Frames Per Second) of each model (Table VIII). The results show that although the GFM and MGM modules are introduced, the full model GMG-L only slightly increases in parameters and FLOPs compared to the baseline, while maintaining an inference speed of 49.11 FPS and a training complexity within an acceptable range. The lightweight version GMG-s achieves excellent accuracy with a fast inference speed of 52.21 FPS, indicating that the proposed modules significantly improve model performance with minimal impact on training and inference efficiency, demonstrating strong practicality and deployment value.
Additionally, we visualize the performance-parameter relationship of the models. Fig. 10 presents a comprehensive comparison of all models’ performance on CIKM2017 and Taxibj, showing that GMG effectively balances model parameters, computational cost, and performance, achieving the best overall performance.
To analyze how the number of GMG Cell layers affects model performance, we conducted experiments comparing configurations with 2 to 5 layers. As shown in Table IX, increasing the number of layers from 2 to 4 led to notable improvements in accuracy, with the 4-layer setup achieving the best results (MSE of 4.0308 and SSIM of 0.9300). However, further increasing to 5 layers resulted in a slight decline in prediction quality, while the computational cost increased significantly (FLOPs reached 0.196T and FPS dropped to 21.17). These results suggest that model depth plays a key role in balancing predictive accuracy and computational efficiency, and highlight the performance trade-offs involved in deeper architectures.
To further verify the effectiveness of the proposed Global Focus Module (GFM), we conduct ablation experiments on ConvRNN-based models using the CIKM2017 and TaxiBJ datasets, with results shown in Table X and Table XI. The experimental results demonstrate that introducing GFM consistently improves performance across both datasets for all ConvRNN variants, significantly reducing prediction errors (MSE and MAE) and enhancing structural similarity (SSIM), without altering the backbone architecture. This clearly shows that GFM, as a plug-and-play module, can stably and effectively enhance the predictive accuracy and robustness of ConvRNN models in different scenarios.
To evaluate the role of the MGM module in MotionRNN, we compare MotionRNN models employing MotionGRU and those incorporating the MGM module on the CIKM2017, Shanghai2020, and TaxiBJ datasets, with the results shown in Table XII. It can be seen that across all datasets, MotionRNN with the MGM module achieves lower MSE and MAE and higher SSIM, exhibiting consistent and significant performance improvements. Notably, the MGM module is inspired by the modeling of non-rigid deformations, enabling it to effectively capture the growth, dissipation, and deformation of targets during motion. This allows the model to demonstrate stronger adaptability and generalization in complex spatiotemporal dynamics. These results indicate that MGM is effective not only for typical non-stationary sequences such as meteorological and traffic data but also, as a plug-and-play module, holds broader application prospects, further validating the rationality and practical value of its design.
Figure 10: The comparison for model performance and efficiency
TABLE VIII: Ablation study results on the CIKM2017 dataset. A checkmark (✓) indicates the presence of a module, while a cross (✗) indicates its absence. ”Simplified” denotes a lightweight variant of GFM.
Model index GFM SAM MGM Params FLOPs FPS MSE (↓) MAE (↓) SSIM (↑)
PredRNN 23.8M 85.43G 58.63 28.8126 159.404 0.7303
MotionRNN MotionGRU 26.9M 94.80G 59.57 27.2091 155.827 0.7406
(A) 26.9M 94.80G 57.37 26.4196 153.801 0.7462
(B) 31.4M 99.67G 55.65 26.9540 155.525 0.7444
(C) 30.4M 94.80G 56.02 26.7479 155.658 0.7448
(D) 31.4M 0.112T 52.57 25.2383 151.723 0.7528
GMG-s (Ours) Simplified 29.2M 0.104T 52.21 25.3616 151.334 0.7501
GMG-m (Ours) 30.4M 0.107T 55.67 26.1861 152.472 0.7487
GMG-L (Ours) 31.44M 0.112T 49.11 25.0215 149.493 0.7513
TABLE IX: Ablation on GMG Layer Number on Shanghai2020.
Layer Params FLOPs FPS MSE MAE SSIM
2 13.67M 66.524G 65.59 5.8221 44.498 0.9033
3 22.58M 0.110T 38.57 4.9531 39.828 0.9165
4 31.44M 0.152T 27.54 4.0308 35.771 0.9300
5 40.41M 0.196T 21.17 4.3757 36.300 0.9260
TABLE X: Ablation study of the Global Focus Module (GFM) on CIKM2017 dataset.
Model MSE (↓) MAE (↓) SSIM (↑)
ConvLSTM 28.8162 161.242 0.7359
+ GFM 26.6154 ($\Delta$-2.2008) 157.141 ($\Delta$-4.1010) 0.7436 ($\Delta$+0.0077)
PredRNN 28.8126 159.404 0.7303
+ GFM 26.7479 ($\Delta$-2.0647) 155.658 ($\Delta$-3.7460) 0.7448 ($\Delta$+0.0145)
PredRNNpp 28.0987 159.258 0.7362
+ GFM 26.9204 ($\Delta$-1.1783) 156.211 ($\Delta$-3.0470) 0.7438 ($\Delta$+0.0076)
MotionRNN 27.2091 155.827 0.7406
+ GFM 26.1861 ($\Delta$-1.0230) 152.472 ($\Delta$-3.3550) 0.7487 ($\Delta$+0.0081)
PredRNNv2 28.0085 161.508 0.7374
+ GFM 27.3749 ($\Delta$-0.6336) 160.649 ($\Delta$-0.8590) 0.7401 ($\Delta$+0.0027)
TABLE XI: Ablation study of the Global Focus Module (GFM) on TaxiBJ dataset.